Using the following six digits, 1, 2, 4, 6, 8, and 9, fill in the blanks below to make the maximum product that you can. Use each digit only once.
__ __ __ x __ __ __

Answer by Edwin McCravy(20060)


It's obvious that the largest two digits should go 
first in each number.

8 _ _ * 9 _ _

And it's obvious that the next two largest 
6 and 4 should go next. But there are two
possibilities for that:

8 6 _ * 9 4 _  or 8 4 _ * 9 6 _

So we just have four possibilities to try:

861 * 942 = 811062
862 * 941 = 811142
841 * 962 = 809042  
842 * 961 = 809162

So the winner is:

862*941 = 811142



In case you want to know the smallest product it's 148*269 = 39812

Edwin